OpenStudy (seratul):

Solve the system of equations: 2x+y-z=-5 4x-2y+z=10 2x+3y+2z=3

OpenStudy (seratul):

@johnweldon1993

OpenStudy (johnweldon1993):

Any ideas how to start?

OpenStudy (seratul):

Um, I guess there are multiple ways to start. After doing this problem a million times, I still couldn't get the third equation to agree with the rest. But I guess we can isolate the z. So if we take the first equation and isolate the z, it would be z=2x+y+5

OpenStudy (johnweldon1993):

Oh good you already worked it! Alright let's go through it and see what went off :) Good first step...next?

OpenStudy (seratul):

Okay, then we would plug it into the second equation right? 4x-2y+2x+y+5=10

OpenStudy (seratul):

6x-y=5 Good so far?

OpenStudy (johnweldon1993):

Looks good

OpenStudy (seratul):

Um, we can isolate the y. y=6x-5

OpenStudy (seratul):

I don't know what would be the best thing to do from here.

OpenStudy (johnweldon1993):

Okay from there *NOT isolating the 'y' * we don't want to further manipulate that equation we have \[\large 6x - y = 5\] From here...take your 'z' you isolated using the first equation....and now plug that in for the 'z' ALSO in the third equation

OpenStudy (johnweldon1993):

Because in a system of 3 equations...your first goal should be to get it down to 2 equations and 2 unknowns

OpenStudy (seratul):

Oh, i see.

OpenStudy (seratul):

Alright. 2x+3y+2(2x+y+5)=3 6x+5y=7

OpenStudy (seratul):

Oh, we can do 7-5y=5+y right?

OpenStudy (johnweldon1993):

Re-check that real quick :)

OpenStudy (seratul):

-7, my bad.

OpenStudy (seratul):

Would it work if I do -7-5y=5+y Cause they both have 6x.

OpenStudy (johnweldon1993):

There we go...so we are down to 2 equations \[\large 6x - y = 5\] \[\large 6x + 5y = -7\] And you're right that WOULD work! So what do you get for 'y' then?

OpenStudy (seratul):

Wait, let me write this down :P

OpenStudy (seratul):

-6y=12 y=-2

OpenStudy (johnweldon1993):

Awesome! Now use that to get your 'x'

OpenStudy (seratul):

Do I plug it into all the equations?

OpenStudy (johnweldon1993):

Not yet...so we went from 3 equations to 2 equations to 1 --> we solved for 'y' Now work backwards, back to the 2 equations, and plug in that 'y' and solve for 'x'

OpenStudy (johnweldon1993):

Reason being, if you plug in 'y' into the original 3 equations, you'll have too much to do to simplify...easier to just work backwards now :)

OpenStudy (seratul):

Oh, okay. So umm, 4x-2(-2)+2x+-2+5=10 4x+4+2x-2+5=10 6x+7=10 6x=3 x=.5

OpenStudy (seratul):

And z should be 4.

OpenStudy (johnweldon1993):

Haha you could have worked with the simplified ones :) \[\large 6x-y=5\] \[\large 6x + 5y = -7\] Regardless, correct...x = 1/2 NOW you have both ... nvm beat me to it...it is indeed lol

OpenStudy (johnweldon1993):

There you go!

OpenStudy (seratul):

Thank you very much. I spent all my free periods trying to finish this up. I appreciate it!

OpenStudy (johnweldon1993):

Not a problem!